# MedTransformer: Accurate Alzheimer’s Disease Diagnosis for 3D MRI Images through 2D Vision Transformers

### 12 Jan 2024 v1

@@ 4.5 Visualization Result
We visualize the activated area our model focusing on based on the transformer attention map. Figure 4 shows a NC-related attention map in 3D MRI images from ADNI dataset in axial, sagittal and coronal views. Because our transformer model has 4 special encoders, we visualize the attention result after each encoder.
Figure 4: Attention map for Normal Control result.
Figure 5: Attention map for Alzheimer’s Disease result.
The attention mostly focuses on the areas that are without the brain mass (thus the atrophic areas of the brain), connecting to the recommended diagnosis procedure of Alzheimer’s from MRI, based on the understanding of its pathology [17].
@@ 5 Conclusions
We proposed a 3D medical image classification method, called MedTransformer, that uses various 2D transformer encoder blocks for Alzheimer’s disease diagnosis. The proposed method uses shared self-attention encoders across different view dimensions, dimension-specific self-attention encoders, intra-dimension cross-attention encoders, and inter-dimension cross-attention encoders to extract and combine information from high-dimensional 3D MR images, with novel techniques such as fusion attention mechanism and morphology augmentation. The experiments show that MedTransformer can achieve outstanding performance compared to various 3D image classification networks in multi-institutional test datasets. The visualization results show that MedTransformer can successfully focus on AD-related regions of 3D MRI images, guiding clinical research on Alzheimer’s Disease.

### 29 Feb 2024 v2

@@ 4.5 Visualization Result
We visualize the activated area our model focusing on based on the transformer attention map. Figure 4 shows a NC-related attention map in 3D MRI images from ADNI dataset in sagittal, coronal and axial views. Because ADAPT has 4 special encoders, we visualize the attention result after each encoder.
Figure 4: Attention map for Normal Control result. Each line corresponds to one view dimension: saggital, coronal and axial.
Figure 5: Attention map for Alzheimer’s Disease result. Each line corresponds to one view dimension: saggital, coronal and axial.
We found that for NC and AD result, the attention mostly focused on some special brain tissues, such as hippocampus, cortex, ventricle and frontal lobe. Disruption of the frontal lobes and its associated networks are a common consequence of neurodegenerative disorders (Sawyer et al., 2017), as well as the hippocampus is most notably damaged by AD (Xu et al., 2021). Based on these understandings of Alzheimer’s pathology (Frisoni et al., 2010), ADAPT successfully captured the AD-related part because with the procedure of Alzheimer’s, the hippocampus and cortex begin to atrophy, and the ventricle begins to expand, which can serve as an evidence of morphology augmentation and confirm the reliability of our proposed ADAPT.
@@ 5 Conclusions
We proposed a 3D medical image classification model, called ADAPT, that uses various 2D transformer encoder blocks for Alzheimer’s disease diagnosis. The proposed method uses shared self-attention encoders across different view dimensions, dimension-specific self-attention encoders, intra-dimension cross-attention encoders, and inter-dimension cross-attention encoders to extract and combine information from high-dimensional 3D MRI images, with novel techniques such as fusion attention mechanism and morphology augmentation. With different encoders, our adaptive training strategy can allow physicians to pay more attention to different dimensions of MRI images. The experiments show that ADAPT can achieve outstanding performance while utilizing the least memory compared to various 3D image classification networks in multi-institutional test datasets. The visualization results show that ADAPT can successfully focus on AD-related regions of 3D MRI images, guiding accurate and efficient clinical research on Alzheimer’s Disease.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning and Alzheimer’s Diagnosis. We hope our work can provide insights for medical doctors and help them diagnose AD in a convenient way. We suggest that our model should be used under human supervision to ensure a perfect result. There are many other potential societal consequences of our work, none which we feel must be specifically highlighted here.